as part of flight training, a student pilot was required to fly to an airport and then return. the average speed on the way to the airport was 115mph, and the average speed returning was 172.5 mph. find the distance between the two airports if the total flying time was 5 h.
Answer by mananth(16946)

let distance be x km
time going + time return = 5 hours
t=d/r
x/115 + x/172.5 = 5
172 1/2= 345/2
x/115 + 2x/345= 5
3x+2x=5*345=1725
5x=1725
x= 345 km